# Efficient Transformed Gaussian Process State-Space Models for Non-Stationary High-Dimensional Dynamical Systems

### v3

@@ III-C EnKF-aided ELBO Evaluation
This section describes how to evaluate $p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})$ in Eq. (26) using the EnKF. First, note that
$p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})$ $=\int\underbracket{p({\mathbf{y}}_{t}|{\mathbf{x}}_{t})}_{\text{likelihood}}\underbracket{p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})}_{\text{prediction}}\mathrm{d}{\mathbf{x}}_{t},$ (27)
@@
Consequently, we employ the EnKF, a widely used method for handling nonlinearities and high-dimensional state spaces [50], to estimate the filtering distribution $p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t})$ at each time step $t$. The details are as follows.
@@ III-C1 Prediction
Given the ensemble of states $\{{\mathbf{x}}_{t-1}^{(i)}\}_{i=1}^{N}$ from the posterior distribution $p({\mathbf{x}}_{t-1}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t-1})$ at time $t-1$, and conditioning on $\mathbf{w}\sim q(\mathbf{w})$ and $\tilde{f}\sim q(\tilde{f})$, we use the state transition in Eq. (29) to generate $N$ predictive samples:
@@ III-C2 Update
The prediction and update steps are fully differentiable, as the sampled latent states become differentiable with respect to the state transition function $\bm{f}(\cdot)$ and the noise covariances $\mathbf{Q}$ and $\mathbf{R}$. This allows for end-to-end gradient-based optimization (e.g.using Adam [51], as described in Algorithm 1), making the framework practical for joint learning and inference [52, 37].
Based on the outlined prediction and update steps, the likelihood $p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})$ in Eq. (27) can be easily evaluated. At each time step $t$, we have:
@@
We now proceed to evaluate the variational lower bound $\mathcal{L}$, as defined in Eq. (26). At each iteration, we apply the reparameterization trick (see e.g. Eq. (33b)) to sample from the Gaussian variational distributions $q(\mathbf{w})$ and $q(\tilde{f})$, enabling unbiased Monte Carlo estimates of the expected log-likelihood term $\mathbb{E}_{q(\mathbf{w},\tilde{f})}[\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})]$. Thanks to the reparameterization trick [53], the ELBO $\mathcal{L}$ remains differentiable with respect to both the model parameters $\bm{\vartheta}=\{\bm{\theta}_{gp},\bm{\psi},\mathbf{Q},\mathbf{R}\}$ and the variational parameters ${\bm{\zeta}}=\{\mathbf{m}_{0},\mathbf{L}_{0},\mathbf{m}_{\mathrm{w}},\bm{\sigma}_{\mathrm{w}}^{2},\mathbf{m},\mathbf{S},\mathbf{Z}\}$. We leverage modern automatic differentiation tools such as PyTorch to compute gradients via backpropagation through time (BPTT), and apply gradient-based optimization methods (e.g., Adam [51]) to maximize $\mathcal{L}$ [54]. The complete procedure for training ETGPSSM with EnKF-based latent state inference is summarized in Algorithm 1.
@@
5:         Get $\tilde{f}$ from $q(\tilde{f})$ using Eq. (8) with ${\mathbf{x}}_{*}={\mathbf{x}}_{t-1}$
@@
$L_{\ell}\leftarrow L_{\ell}+\log p({\mathbf{y}}_{t}|{\mathbf{y}}_{1:t-1},\mathbf{w},\tilde{f})\vspace{-.1in}$
@@
15:     Maximize $\mathcal{L}$ and update $\bm{\vartheta}$, ${\bm{\zeta}}$ using Adam [51]
@@
Remark 4 (Efficiency).
The proposed ETGPSSM in Algorithm 1 improves upon existing GPSSMs in two key aspects of efficiency. First, it reduces the computational complexity from $\mathcal{O}(d_{x}M^{3})$ to $\mathcal{O}(M^{3})$ by relying on a single GP, while the added cost of a shallow neural network remains negligible. Second, by sharing the GP and flow network across dimensions, it avoids the quadratic scaling of variational parameters associated with independent GPs, as discussed in Section II-C. Furthermore, by replacing explicit variational parameterization of the latent states with EnKF-based inference, the method significantly reduces parametric overhead. Together, these design choices lead to substantial gains in both computational and modeling efficiency. Empirical results validating these advantages are presented in Section IV-B.
In the next section, we empirically evaluate the performance of ETGPSSM across a range of dynamical system modeling tasks.
            Data               AD-EnKF [52] AD-EnKF (bnn)          EnVI [37] ETGPSSM (dnn)      ETGPSSM (bnn)
Figure 4: Non-stationary kink transition function learning performance (mean $\pm$ $2\sigma$) using various methods across different levels of emission noise ($\sigma_{\mathrm{R}}^{2}\in\{0.0008,0.008,0.08,0.8\}$, from top to bottom). The blue curve () represents the learned mean function, while the red line () indicates the true system transition function. The shaded blue region depicts the confidence interval.
@@ IV Experiments
We begin by evaluating ETGPSSM’s ability to model non-stationary dynamical systems in Section IV-A. We then assess its computational and parametric efficiency, as well as filtering accuracy, in chaotic high-dimensional settings in Section IV-B. Lastly, Section IV-C highlights the model’s predictive performance across a range of real-world time series datasets. Unless otherwise stated, all GPSSMs use the squared exponential (SE) kernel [9], with $20$ inducing points initialized at random. The shallow neural network used in ETGPSSM is a fully connected three-layer architecture with ReLU activations and hidden dimensions $128$ and $64$, mapping from $\mathbb{R}^{d_{x}}$ to the flow parameter space $\Theta\!\subseteq\!\mathbb{R}^{2d_{x}}$. Additional implementation details are available in the publicly released code. [^5] https://github.com/zhidilin/gpssmProj/tree/main/high_dim_GPSSM
@@ IV-A Non-Stationary Dynamical System Learning
$f({\mathbf{x}})\!=\!\underbrace{\left[0.8+({\mathbf{x}}+0.2)\left(1-\frac{5}{1+\exp(-2{\mathbf{x}})}\right)\right]}_{\text{``kink'' function}}s({\mathbf{x}})-o({\mathbf{x}})$ (37)
@@
The kink function is a standard benchmark for evaluating GPSSMs [27, 31, 37]. Here, we extend it with non-stationary slope and oscillatory terms to increase modeling complexity (see Fig. 4). We generate training data ${\mathbf{y}}_{1:T}$ by fixing $\sigma_{\mathrm{Q}}^{2}$ at $0.05$ and varying $\sigma_{\mathrm{R}}^{2}$ systematically across the set $\{0.0008,0.008,0.08,0.8\}$, producing four noise regimes. ETGPSSM is compared to two leading baselines: the neural network-based AD-EnKF [52] and the GP-based EnVI [37]. Consistent with previous works [31, 37], all methods share the true emission model while learning the transition function, ensuring a fair comparison. The system dynamics learning results across different methods are shown in Fig. 4, where our method is represented in the last two columns. The suffixes “dnn” and “bnn” in brackets indicate the network types—Bayesian neural network (BNN) or deep neural network (DNN)—used in our model (see Fig. 3).
The results in Fig. 4 show that, across all noise levels, ETGPSSM achieves consistently superior system identification performance. In the low-noise regimes($\sigma_{\mathrm{R}}^{2}=\{0.0008,0.008\}$), the DNN-based ETGPSSM excels at capturing both smooth and oscillatory regions of the transition function. By contrast, AD-EnKF struggles with the smooth left segment and performs well only on the right, highly nonlinear region. EnVI, the strongest GPSSM baseline, fails to capture the non-stationarity and underfits the high-frequency component.
@@ IV-B High-Dimensional Latent State Estimation
In this subsection, we evaluate latent state estimation in the Lorenz-96 system, a standard benchmark for studying chaotic dynamics in weather and climate modeling [55]. The system consists of $d_{x}\in\mathbb{Z}_{+}$ coupled ordinary differential equations (ODEs):
[…]
where $F=8$ induces fully chaotic behavior, with small differences in initial conditions leading to divergent trajectories [55]. The dynamics model is discretized using the Euler method, and at each time step $t$, observations are generated through a linear emission model ($\bm{C}\!=\!\bm{I}$) with Gaussian additive noise, where the noise level is significant with the covariance given by $\mathbf{R}=4\bm{I}$.
We first assess the scalability of ETGPSSM in high-dimensional settings. Fig. 5 compares its parameter growth and runtime to standard GPSSMs, using $100$ inducing points. As shown in the left panel, ETGPSSM exhibits linear growth in parameter count with increasing $d_{x}$, while GPSSM grows quadratically. This is enabled by ETGPSSM’s use of a shared GP combined with normalizing flows, in contrast to GPSSM’s reliance on independent GPs per dimension.
@@
Figure 5: (Left) The number of parameters in ETGPSSM grows linearly with the state dimension $d_{x}$, whereas in GPSSM, it scales quadratically. (Right) ETGPSSM maintains low computational costs as $d_{x}$ increases, while GPSSM exhibits linear growth.
Figure 6: Latent state estimation performance for 100-dimensional Lorenz-96 system, with 9 dimensions uniformly sampled for visualization. Root mean square error (RMSE) values for different methods: 0.5468 for EnKF, 1.2028 for ETGPSSM, 1.6187 for AD-EnKF. RMSE between latent states and observations: 1.9855.
We next fix the dimension $d_{x}\!=\!100$ in the Lorenz-96 system and generate observations of length $T\!=\!600$ for training ETGPSSM and its deep neural network-based competitor, AD-EnKF [52]. The EnVI [37], which already caused out-of-memory errors [^6] All experiments were conducted on a machine equipped with an Intel Xeon W7-3545 v6 processor (24 cores, 2.7 GHz), 128 GB of DDR5 RAM, and an NVIDIA RTX 4500 Ada GPU (24 GB VRAM), running Ubuntu 20.04 LTS. at $d_{x}\!=\!30$ with an ensemble size $N\!=\!150$, is excluded due to its prohibitive computational cost in this high-dimensional setting. The corresponding state inference performance after training is shown in Fig. 6, where the performance of EnKF is also included for comparison.
As depicted in Fig. 6, our ETGPSSM achieves an RMSE of $1.2028$, outperforming AD-EnKF, which yields an RMSE of $1.6187$. Additionally, note that the RMSE of the observations (1.9855) serves as a baseline, demonstrating that ETGPSSM significantly reduces the estimation error compared to the raw observations. This result highlights the superior filtering performance of ETGPSSM in high-dimensional chaotic systems. While EnKF achieves the lowest RMSE of $0.5468$, it assumes full knowledge of the dynamics and performs no learning. In contrast, ETGPSSM simultaneously learns the system and infers latent states, offering a more general and powerful solution for complex, high-dimensional systems.
@@ IV-C Real-World Time Series Prediction
In this subsection, we evaluate the prediction performance of ETGPSSM on several real-world time series datasets [^7] https://homes.esat.kuleuven.be/~smc/daisy/daisydata.html commonly used for system identification (see Table I). In every dataset, the first half of the sequence is utilized as training data, with the remaining portion designated for testing. Standardization of all datasets is conducted based on the training sequence. Table I reports the series prediction results, wherein the RMSE is averaged over 50-step ahead forecasting. We compare ETGPSSM against state-of-the-art methods, including neural network-based approaches—DKF [7] and AD-EnKF [52]—as well as various GPSSMs: vGPSSM [29], CO-GPSSM [36], PRSSM [30], ODGPSSM [39], VCDT [31], and EnVI [37]. Following previous works, the latent space dimension is set to $d_{x}=4$ for all the methods.
TABLE I: The 50-step ahead prediction performance in terms of RMSE of different models (all with $d_{x}=4$) on system identification datasets. Mean and standard deviation of the prediction results are shown as an averaged result over ten different seeds. The three lowest RMSE values are highlighted in shades of green, with a deeper one indicating lower RMSE.
@@
DKF [7] $1.204\pm 0.250$ $0.144\pm 0.005$ $0.735\pm 0.001$ $1.465\pm 0.087$ $5.589\pm 0.066$
AD-EnKF [52] $0.705\pm 0.117$ $0.057\pm 0.006$ $0.756\pm 0.114$ $0.182\pm 0.053$ $1.408\pm 0.090$
AD-EnKF (bnn) $0.705\pm 0.088$ $0.053\pm 0.007$ $0.896\pm 0.088$ $0.155\pm 0.030$ $1.361\pm 0.061$
vGPSSM [29] $1.640\pm 0.011$ $0.268\pm 0.414$ $0.740\pm 0.010$ $0.822\pm 0.002$ $3.676\pm 0.145$
CO-GPSSM [36] $0.803\pm 0.011$ $0.079\pm 0.018$ $0.736\pm 0.007$ $0.366\pm 0.146$ $1.898\pm 0.157$
PRSSM [30] $0.691\pm 0.148$ $0.074\pm 0.010$ $0.647\pm 0.057$ $0.174\pm 0.013$ $1.503\pm 0.196$
ODGPSSM [39] $0.666\pm 0.074$ $0.068\pm 0.006$ $0.708\pm 0.052$ $0.171\pm 0.011$ $1.704\pm 0.560$
VCDT [31] $0.815\pm 0.012$ $0.065\pm 0.005$ $0.735\pm 0.005$ $0.667\pm 0.266$ $2.052\pm 0.163$
EnVI [37] $0.657\pm 0.095$ $0.055\pm 0.002$ $0.703\pm 0.050$ $0.125\pm 0.017$ $1.388\pm 0.123$
ETGPSSM (dnn) $0.646\pm 0.081$ $0.050\pm 0.003$ $0.668\pm 0.092$ $0.137\pm 0.030$ $1.300\pm 0.052$
ETGPSSM (bnn) $0.646\pm 0.095$ $0.053\pm 0.002$ $0.703\pm 0.028$ $0.154\pm 0.026$ $1.313\pm 0.048$
As reported in Table I, ETGPSSM demonstrates strong overall performance. On the Actuator dataset, both ETGPSSM (dnn) and ETGPSSM (bnn) achieve an RMSE of $0.646$, outperforming EnVI ($0.657$) and significantly improving upon AD-EnKFs, which yield an RMSE of $0.705$. Similar trends are observed on the Ball Beam and Gas Furnace datasets, where ETGPSSM ranks first and second, respectively, indicating effective modeling of system dynamics. On the Drive and Dryer datasets, however, GPSSMs such as PRSSM and EnVI attain slightly better performance. This is likely due to the relatively low latent state dimension ($d_{x}=4$), where conventional GPSSMs, including PRSSM, ODGPSSM, and EnVI, remain competitive, as discussed in [37]. Nonetheless, we highlight that although the primary contribution of the proposed ETGPSSM lies in its scalability and efficiency in high-dimensional settings, it also remains highly competitive in low-dimensional scenarios, achieving superior or very comparable performance.

### v4

@@ III-C EnKF-aided ELBO Evaluation
[…]
@@
Consequently, we employ the EnKF, a widely used method for handling nonlinearities and high-dimensional state spaces [evensen1994sequential], to estimate the filtering distribution $p({\mathbf{x}}_{t}|\mathbf{w},\tilde{f},{\mathbf{y}}_{1:t})$ at each time step $t$. The details are as follows.
@@ III-C1 Prediction
[…]
@@ III-C2 Update
The prediction and update steps are fully differentiable, as the sampled latent states become differentiable with respect to the state transition function $\bm{f}(\cdot)$ and the noise covariances $\mathbf{Q}$ and $\mathbf{R}$. This allows for end-to-end gradient-based optimization (e.g.using Adam [kingma2015adam], as described in Algorithm 1), making the framework practical for joint learning and inference [chen2022autodifferentiable, lin2023ensemble].
[…]
@@
We now proceed to evaluate the variational lower bound $\mathcal{L}$, as defined in Eq. (26). At each iteration, we apply the reparameterization trick to sample from the Gaussian variational distributions $q(\mathbf{w})$, $q(\mathbf{x}_{0})$, and q(f~)q($\tilde${f}). For instance, a sample from q(f~)q($\tilde${f}) at input $\mathbf{x}_{t}$ is generated as
$\tilde{f}(\mathbf{x}_{t})=\xi_{p}(\mathbf{x}_{t})+\sqrt{\Xi_{p}(\mathbf{x}_{t})}\cdot\epsilon,\quad\epsilon\sim\mathcal{N}(0,\,1),$
where $\xi_{p}(\cdot)$ and $\Xi_{p}(\cdot)$ are given by Eq. (9). Similarly, for $q(\mathbf{w})$ and $q(\mathbf{x}_{0})$, we also use the standard Gaussian reparameterization. This enables unbiased Monte Carlo estimates of the expected log-likelihood term in $\mathcal{L}$. And thanks to the reparameterization trick [kingma2019introduction], the ELBO $\mathcal{L}$ remains differentiable with respect to both the model parameters $\bm{\vartheta}\!=\!\{\bm{\theta}_{gp},\bm{\psi},\mathbf{Q},\mathbf{R}\}$ and the variational parameters ${\bm{\zeta}}\!=\!\{\mathbf{m}_{0},\mathbf{L}_{0},\mathbf{m}_{\mathrm{w}},\bm{\sigma}_{\mathrm{w}}^{2},\mathbf{m},\mathbf{S},\mathbf{Z}\}$. We leverage modern automatic differentiation tools such as PyTorch to compute gradients via backpropagation through time, and apply gradient-based optimization methods (e.g., Adam [kingma2015adam]) to maximize $\mathcal{L}$ [paszke2019pytorch]. The complete procedure for training ETGPSSM with EnKF-based latent state inference is summarized in Algorithm 1.
@@
5:   Get $\tilde{f}({\mathbf{x}}_{t-1}^{(i)})$ using Eq. (8) with ${\mathbf{x}}_{*}={\mathbf{x}}_{t-1}^{(i)},\forall i$
@@
[…]
@@
15:  Maximize $\mathcal{L}$ and update $\bm{\vartheta}$, $\bm{\zeta}$} using Adam [kingma2015adam]
@@
Remark 4.
The proposed ETGPSSM in Algorithm 1 improves upon existing GPSSMs in several key aspects of efficiency. First, it reduces the computational complexity from $\mathcal{O}(d_{x}M^{3})$ to $\mathcal{O}(M^{3})$ by relying on a single GP, while the added cost of a shallow neural network remains negligible. Second, by sharing the GP and flow network across dimensions, it avoids the quadratic scaling of variational parameters associated with independent GPs, as discussed in Section II-C. Furthermore, by replacing explicit variational parameterization of the latent states with EnKF-based inference, the method significantly reduces parametric overhead. Together, these design choices lead to substantial gain in both computational and modeling efficiency. The overall computational complexity is $\mathcal{O}(M^{3}+d_{x}^{3})$, where the $\mathcal{O}(d_{x}^{3})$ term comes from the EnKF operations.
It is important to note that in practice, $M$ is typically selected based on the complexity of the transition dynamics and the available computational budget, resulting in a trade-off between computation and modeling accuracy. Nevertheless, further investigation into how $M$ should scale with $d_{x}$ to preserve approximation accuracy in high-dimensional GPSSMs [burt2020convergence] is warranted for future research.
      Data       AD-EnKF [chen2022autodifferentiable] AD-EnKF (bnn)     EnVI [lin2023ensemble] ETGPSSM (dnn)    ETGPSSM (bnn)
Figure 4: Non-stationary kink transition function learning performance (mean $\pm$ $2\sigma$) using various methods across different levels of emission noise ($\sigma_{\mathrm{R}}^{2}\in\{0.0008,0.008,0.08,0.8\}$, from top to bottom). The blue curve () represents the learned mean function, while the red line () indicates the true system transition function. The shaded region depicts the total predictive uncertainty for the state transition, which includes the estimated transition noise level $(\pm 2\,\hat{\sigma}_{Q})$ for all methods, plus function uncertainty for probabilistic models (BNN/GP).
@@ IV Experiments
In this section, we empirically evaluate the performance of ETGPSSM across a range of tasks. We begin by evaluating ETGPSSM’s ability to model non-stationary dynamical systems in Section IV-A. We then assess its computational and parametric efficiency, as well as filtering accuracy, in chaotic high-dimensional settings in Section IV-B. Lastly, Section IV-C highlights the model’s predictive performance across a range of real-world time series datasets. Unless otherwise stated, all GPSSMs use the SE kernel [williams2006gaussian], with $20$ inducing points initialized at random. The shallow neural network used in ETGPSSM is a fully connected three-layer architecture with ReLU activations and hidden dimensions $128$ and $64$, mapping from $\mathbb{R}^{d_{x}}$ to the flow parameter space $\Theta\!\subseteq\!\mathbb{R}^{2d_{x}}$. For further implementation details, we refer the reader to the supplementary material and the public code repository. [^4] https://github.com/zhidilin/gpssmProj/tree/main/high_dim_GPSSM
@@ IV-A Non-Stationary Dynamical System Learning
[…]
@@
The kink function is a standard benchmark for evaluating GPSSMs [frigola2014variational, ialongo2019overcoming, lin2023ensemble], and is readily modeled by a stationary GP with an SE kernel. Here, we break this stationarity by extending it with input-dependent slope and oscillatory terms, creating a non-stationary benchmark that presents a greater challenge for dynamical models (see Fig. 4). We generate training data ${\mathbf{y}}_{1:T}$ by fixing $\sigma_{\mathrm{Q}}^{2}$ at $0.05$ and varying $\sigma_{\mathrm{R}}^{2}$ systematically across the set $\{0.0008,0.008,0.08,0.8\}$, producing four noise regimes. ETGPSSM is compared to two leading baselines: the neural network-based AD-EnKF [chen2022autodifferentiable] and the GP-based EnVI [lin2023ensemble]. The AD-EnKF employs a fully-connected network with ReLU activations and hidden layers of dimensions 128 and 64, to model the state transition. Further implementation details (e.g., learning rate, optimizer) are provided in Supplement B. Consistent with previous works [ialongo2019overcoming, lin2023ensemble], all methods share the true emission model while learning the transition function, ensuring a fair comparison. The system dynamics learning results across different methods are shown in Fig. 4, where our method is represented in the last two columns. The suffixes “dnn” and “bnn” in brackets indicate the network types—Bayesian neural network (BNN) or deep neural network (DNN)—used in our model (see Fig. 3).
Figure 5: (Left) Comparison of the total number of parameters (including variational and model parameters). In ETGPSSM, the count is dominated by the GP ($Md_{x}+3+M+M^{2}$) and the neural network with architecture $d_{x}\!\to\!128\!\to\!64\!\to\!2d_{x}$ ($258\,d_{x}+8384$), resulting in overall linear scaling in $d_{x}$. In GPSSM, the parameters are dominated by the $d_{x}$ independent GPs, ($Md_{x}^{2}+(3+M+M^{2})d_{x}$), leading to quadratic scaling in $d_{x}$. (Right) ETGPSSM maintains low computational costs as $d_{x}$ increases, while GPSSM exhibits linear growth.
The results in Fig. 4 show that, across all noise levels, ETGPSSM achieves consistently superior system identification performance. In the low-noise regimes($\sigma_{\mathrm{R}}^{2}=\{0.0008,0.008\}$), the DNN-based ETGPSSM excels at capturing both smooth and oscillatory regions of the transition function. By contrast, AD-EnKF struggles with the smooth left segment and performs well only on the right, highly nonlinear region. EnVI, the state-of-the-art GPSSM baseline, fails to capture the non-stationarity and underfits the high-frequency component.
@@ IV-B High-Dimensional Latent State Estimation
In this subsection, we evaluate latent state estimation in the Lorenz-96 system, a standard benchmark for studying chaotic dynamics in weather and climate modeling [lorenz1996predictability]. The system consists of $d_{x}\in\mathbb{Z}_{+}$ coupled ordinary differential equations (ODEs):
$\frac{d\mathrm{x}_{d}}{dt}=(\mathrm{x}_{d+1}-\mathrm{x}_{d-2})\mathrm{x}_{d-1}-\mathrm{x}_{d}+F,\ 1\leq d\leq d_{x}$ (40)
where $F=8$ induces fully chaotic behavior, with small differences in initial conditions leading to divergent trajectories [lorenz1996predictability]. The dynamics model is discretized using the Euler method, and at each time step $t$, observations are generated through a linear emission model ($\bm{C}\!=\!\bm{I}$) with Gaussian additive noise, where the noise level is significant with the covariance given by $\mathbf{R}=4\bm{I}$.
We first assess the scalability of ETGPSSM in high-dimensional settings. For simplicity, we focus exclusively on the dnn variant of ETGPSSM. The bnn variant achieves comparable performance but doubles the number of parameters by introducing variance parameters for each neural network weight. This observation aligns with prior work showing that, on sufficiently large datasets, the bnn variant offers no significant performance gain over its deterministic counterpart [maronas2022efficient, maronas2021transforming]. Figure 5 shows the parameter count and runtime of ETGPSSM and standard GPSSMs, both using 100 inducing points. As shown in the left panel, ETGPSSM exhibits linear growth in parameter count with increasing $d_{x}$, while GPSSM grows quadratically. This is enabled by ETGPSSM’s use of a shared GP combined with normalizing flows, in contrast to GPSSM’s reliance on independent GPs per dimension.
@@
Figure 6: Latent state estimation performance for a 100-dimensional Lorenz-96 system, with 3 dimensions sampled for visualization. Root mean square error (RMSE) values for different methods are 0.4468 for EnKF, 1.1583 for ETGPSSM, and 1.4553 for AD-EnKF. RMSE between the true latent states and the observations is 1.9855, which serves as a baseline, indicating the error if observations were used directly as state estimates.
TABLE I: Comparison of filtering performance in the pendulum system.
Method rmse ($\downarrow$) coverage spread crps ($\downarrow$) ad-enkf 1.4553 0.3007 0.2885 0.2681 etgpssm 1.1583 0.3587 0.2774 0.2101 enkf 0.4468 0.5866 0.1741 0.0685
We next fix the dimension $d_{x}\!=\!100$ in the Lorenz-96 system and generate observations of length T=600T\!=\!600 for training ETGPSSM and its deep neural network-based competitor, AD-EnKF [chen2022autodifferentiable]. The EnVI [lin2023ensemble], which already caused out-of-memory errors [^5] All experiments were conducted on a machine equipped with an Intel Xeon W7-3545 v6 processor (24 cores, 2.7 GHz), 128 GB of DDR5 RAM, and an NVIDIA RTX 4500 Ada GPU (24 GB VRAM), running Ubuntu 20.04 LTS. at $d_{x}\!=\!30$ with an ensemble size N=150N\!=\!150, is excluded due to its prohibitive computational cost in this high-dimensional setting. The corresponding state inference performance after training is summarized in Fig. 6 and Table I. Table I presents four standard filtering metrics: RMSE, spread, coverage probability, and CRPS [spantini2022coupling, gneiting2007probabilistic]. Detailed definitions of these metrics are provided in Supplement B. For comparative analysis, the performance of EnKF is also included.
As shown in Fig. 6 and Table I, ETGPSSM achieves superior performance compared to AD-EnKF in both mean estimation (see e.g. RMSE) and uncertainty quantification (see e.g. CRPS). Notably, with the observation RMSE of 1.9855 serving as a baseline, ETGPSSM demonstrates substantial error reduction relative to the raw observational data. This result highlights the superior filtering performance of ETGPSSM in high-dimensional chaotic systems. While EnKF achieves the best filtering performance, e.g., lowest RMSE, it assumes full knowledge of the dynamics and performs no learning. In contrast, ETGPSSM simultaneously learns the system and infers latent states, offering a more general and powerful solution for complex, high-dimensional systems.
@@ IV-C Real-World Time Series Prediction
In this subsection, we evaluate the prediction performance of ETGPSSM on several real-world time series datasets [^6] https://homes.esat.kuleuven.be/~smc/daisy/daisydata.html commonly used for system identification (see Table II). In every dataset, the first half of the sequence is utilized as training data, with the remaining portion designated for testing. Standardization of all datasets is conducted based on the training sequence. Table II reports the series prediction results, wherein the RMSE is averaged over 50-step ahead forecasting. We compare ETGPSSM against state-of-the-art methods, including neural network-based approaches—DKF [krishnan2017structured] and AD-EnKF [chen2022autodifferentiable]—as well as various GPSSMs: vGPSSM [eleftheriadis2017identification], CO-GPSSM [lin2023towards], PRSSM [doerr2018probabilistic], ODGPSSM [lin2022output], VCDT [ialongo2019overcoming], and EnVI [lin2023ensemble]. Following previous works, the latent space dimension is set to $d_{x}=4$ for all the methods.
TABLE II: The 50-step ahead prediction performance in terms of RMSE of different models (all with $d_{x}=4$) on system identification datasets. Mean and standard deviation of the prediction results are shown as an averaged result over ten different seeds. The three lowest RMSE values are highlighted in shades of green, with a deeper one indicating lower RMSE.
@@
DKF [krishnan2017structured] $1.204\pm 0.250$ $0.144\pm 0.005$ $0.735\pm 0.001$ $1.465\pm 0.087$ $5.589\pm 0.066$
AD-EnKF [chen2022autodifferentiable] $705\pm 0.117 0.0$$057\pm 0.006$ $0.756\pm 0.114$ $0.182\pm 0.053$ $1.408\pm 0.090$
AD-EnKF (bnn) $705\pm 0.088 0.0$$053\pm 0.007$ $0.896\pm 0.088$ $0.155\pm 0.030$ $1.361\pm 0.061$
vGPSSM [eleftheriadis2017identification] $1.640\pm 0.011$ $0.268\pm 0.414$ $0.740\pm 0.010$ $0.822\pm 0.002$ $3.676\pm 0.145$
CO-GPSSM [lin2023towards] $803\pm 0.011 0.0$$079\pm 0.018$ $0.736\pm 0.007$ $0.366\pm 0.146$ $1.898\pm 0.157$
PRSSM [doerr2018probabilistic] $691\pm 0.148 0.0$$074\pm 0.010$ $0.647\pm 0.057$ $0.174\pm 0.013$ $1.503\pm 0.196$
ODGPSSM [lin2022output] $666\pm 0.074 0.0$$068\pm 0.006$ $0.708\pm 0.052$ $0.171\pm 0.011$ $1.704\pm 0.560$
VCDT [ialongo2019overcoming] $815\pm 0.012 0.0$$065\pm 0.005$ $0.735\pm 0.005$ $0.667\pm 0.266$ $2.052\pm 0.163$
EnVI [lin2023ensemble] $657\pm 0.095 0.0$$055\pm 0.002$ $0.703\pm 0.050$ $0.125\pm 0.017$ $1.388\pm 0.123$
ETGPSSM (dnn) $646\pm 0.081 0.0$$050\pm 0.003$ $0.668\pm 0.092$ $0.137\pm 0.030$ $1.300\pm 0.052$
ETGPSSM (bnn) $646\pm 0.095 0.0$$053\pm 0.002$ $0.703\pm 0.028$ $0.154\pm 0.026$ $1.313\pm 0.048$
As reported in Table II, ETGPSSM demonstrates strong overall performance. On the Actuator dataset, both ETGPSSM (dnn) and ETGPSSM (bnn) achieve an RMSE of $0.646$, outperforming EnVI ($0.657$) and significantly improving upon AD-EnKFs, which yield an RMSE of $0.705$. Similar trends are observed on the Ball Beam and Gas Furnace datasets, where ETGPSSM ranks first and second, respectively, indicating effective modeling of system dynamics. On the Drive and Dryer datasets, however, GPSSMs such as PRSSM and EnVI attain slightly better performance. This is likely due to the relatively low latent state dimension ($d_{x}=4$), where conventional GPSSMs, including PRSSM, ODGPSSM, and EnVI, remain competitive, as discussed in [lin2023ensemble]. Nonetheless, we highlight that although the primary contribution of the proposed ETGPSSM lies in its scalability and efficiency in high-dimensional settings, it also remains highly competitive in low-dimensional scenarios, achieving superior or very comparable performance.